OpenStudy (anonymous):

12-4/5(x+15)=4 Wth??

OpenStudy (xmistermayhem):

\[12-\frac{ 4 }{ 5 }(x+15)=4 \] Is this your question?

OpenStudy (anonymous):

@xmistermayhem yes

OpenStudy (xmistermayhem):

Okay, so this is easy. =) You know how to get x alone on one side?

OpenStudy (anonymous):

I think. You have to multiply 4/5 by X and 15 right?

OpenStudy (xmistermayhem):

Start putting bits and pieces on the other side - Since 12 is positive, and you want to cancel it out, subtract 12 from both sides. \[(12 - \frac{ 4 }{ 5 }(x+15) ) - 12 = (4) -12\]

OpenStudy (xmistermayhem):

Then you get rid of that 5. \[5*(\frac{ 4 }{ 5 }(x+15)) = (-8)*5\]

OpenStudy (xmistermayhem):

What's left is \[4(x+15) = -40\]

OpenStudy (xmistermayhem):

Divide out that 4... \[\frac{ 4(x+15) }{ 5 } = \frac{ -40 }{ 4 }\]

OpenStudy (xmistermayhem):

Left with \[x + 15 = -10\] So get rid of the 15. \[(x + 15) -15 = -10 -15\] Then the answer to that is x.

OpenStudy (anonymous):

Okay thanks! @xmistermayhem

OpenStudy (xmistermayhem):

You're quite welcome. Good luck!